yellow light emitted from a sodium lamp has a wavelength of 580nm.Calculate the frequency and wavenumber of the yellow light

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Yellow light emitted from a sodium lamp has a wavelength (.) of 580 nmwavelength=580nm =580*10^-9mwe havefrequency =velocity of light /wavelength=3*10^8/580*10^-9=5.17*10^14 per secondwavenumber =1/wavelegth=1/580*10^-9=1.724*10^6 m^-1
 
Regards
Arun